find the middle term (2c - 12š¯‘‘)10


1
Expert's answer
2022-04-06T16:35:56-0400

"\left( 2c-12d \right) ^{10}\\The\,\,middle\,\,term\,\,is\,\,for\,\,5th\,\,powers:\\C_{10}^{5}\left( 2c \right) ^5\left( -12d \right) ^5=-2006581248c^5d^5\\"